A division with total assets of £200,000 currently earns a ROI of 10%. It can make an additional investment of £20,000 for a 5 year life with nil residual value. The average operating income per annum from this investment would be £4,000. The division’s cost of capital is 11%.

Required

Compute and comment on the Return on Investment and the Residual Income, with and without the additional investment, and then advise the company whether it should make the additional investment.          

Critique the relative advantages of ROI and RI for measuring divisional performance.
Answer #1

Division’s total assets of £200,000 currently earns a ROI of 10%. It can make an additional investment of £20,000 for a 5 year life with nil residual value.
Cost of Capital =11%
Note- Every Quantity in £

First part:
Without new investment
ROI=10%, which is less than cost of capital
Return is 20000
Residual income= 20000- (11% of 200000) = -2000 which is negative
With new investment
Return= 20000+4000=24000
ROI= 24,000/ (200,000+20,000)= 10.9%
Residual Income= 24000- (11% of 220000)= -200
Since both ROI and Residual income are increasing by the new investment, so Division should go ahead with the investment

Second part-
Benefits of Using RI over ROI=
1. We can see in the above example that even though the ROI is still less than 11% but the residual value has become almost zero from high negative value. So, a new investment might add to residual income which might decrease the ROI also.
2.So, for measuring the performance, the decision can be made for the whole division as a whole instead of just a small unit.
3. Manager can compare cost of units under them in a different way and they can be aware of the changes each unit is making under them
Benefits of using ROI
1. RI is based on size of divisions so one cannot compare two different units of different sizes, ROI is better
2. RI is based on profit measuring in accounting , which is subject to manipulation and ROI is less subject to it.